VIBRATIONAL-ROTATIONAL-TUNNELING SPECTROSCOPY OF N_{2}-D_{2}O WITH AN ALL SOLID STATE, OPTICAL HETERODYNE, SUBMILLIMETER-WAVE SPECTROMETER

underlinePIN CHEN, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; GEOFFREY A. BLAKE, Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125.

Recent advances in the fabrication of ultrafast low-temperature-grown GaAs optoelectronic devices have made possible the generation of spectroscopically significant amounts of tunable, cw far-IR radiation via optical heterodyne processes. We will present VRT spectra of the N_{2}-D_{2}O complex using an all-solid-state, passively frequency-stabilized, optical-heterodyne FIR spectrometer.